# M2D2: Exploring General-purpose Audio-Language Representations Beyond CLAP

### v1

@@ Abstract
Contrastive language-audio pre-training (CLAP) has addressed audio-language tasks such as audio-text retrieval by aligning audio and text in a common feature space. While CLAP addresses general audio-language tasks, its audio features do not generalize well in audio tasks. In contrast, self-supervised learning (SSL) models learn general-purpose audio features that perform well in diverse audio tasks. We pursue representation learning that can be widely used in audio applications and hypothesize that a method that learns both general audio features and CLAP features should achieve our goal, which we call a general-purpose audio-language representation. To implement our hypothesis, we propose M2D2, a second-generation masked modeling duo (M2D) that combines an SSL M2D and CLAP. M2D2 learns two types of features using two modalities (audio and text) in a two-stage training process. It also utilizes advanced LLM-based sentence embeddings in CLAP training for powerful semantic supervision. In the first stage, M2D2 learns generalizable audio features from M2D and CLAP, where CLAP aligns the features with the fine LLM-based semantic embeddings. In the second stage, it learns CLAP features using the audio features learned from the LLM-based embeddings. Through these pre-training stages, M2D2 should enhance generalizability and performance in its audio and CLAP features. Experiments validated that M2D2 achieves effective general-purpose audio-language representation, highlighted with SOTA fine-tuning mAP of 49.0 for AudioSet, SOTA performance in music tasks, and top-level performance in audio-language tasks.
@@
Contrastive language-image pre-training (CLIP)[1] has enabled the alignment of two modalities and has significantly impacted diverse domains. In the audio domain, contrastive language-audio pre-training (CLAP) has extended CLIP to audio[2][3][4][5], followed by various advancements[6][7][8][9][10][11] that have exhibited superior performance for applications such as zero-shot classification and audio captioning[7][12][13]. CLAP can map audio and captions into features aligned in common feature space; therefore, these features can be effective for various downstream application tasks since many task settings are typically related to language, such as inference of class labels.
While the CLAP models have exhibited versatility in audio-language tasks, their audio features do not generalize well in conventional audio tasks because they typically use audio encoders based on supervised learning models. The state-of-the-art (SOTA) audio representations are self-supervised learning (SSL) models known to generalize to diverse downstream tasks[14]. In contrast, the supervised learning models, and thus the CLAP models, suffer on tasks different from the pre-training, as we reported in [11][15]. For example, they perform well in ESC-50[16], an environmental sound classification similar to the pre-trained AudioSet[17] but poorly in VoxCeleb1[18], a speaker identification task. However, some studies have applied CLAP audio features to a task where SSL audio features may perform well, such as [19][20].
@@
On the other hand, to learn the best of CLAP, a SOTA sentence embedding model based on a large-language model (LLM) would be a good candidate as a CLAP text encoder. In the natural language processing (NLP) domain, sentence embedding models have recently progressed to exploit an LLM[21]. LLMs have demonstrated overwhelming ability in language-related tasks[22] and have the potential to map audio captions into better semantic features based on knowledge accumulated in huge parameters learned from massive data.
To implement a general-purpose audio-language representation and achieve its highest potential, we propose M2D2, which extends a SOTA masked prediction-based SSL, M2D[23], combines it with a CLAP objective, and exploits the powerful LLM-based sentence embeddings, as shown in Fig. 1. It uses a two-stage pre-training strategy that suppresses required computational resources, thus enabling multitasking of M2D and CLAP, contrastive learning with larger batch size, and even learning from LLM-based sentence embeddings.
The two-stage pre-training strategy enables learning the best features while avoiding substantial computational costs. In the first stage, M2D2 learns from a multitask of M2D and CLAP and uses LLM-based text features for strong semantic supervision, aiming to learn the best general-purpose audio features that can also be projected to an effective CLAP feature. In the second stage, M2D2 pre-trains a non-LLM text encoder to learn effective CLAP text features that align better with the learned audio features. We aim to exploit the LLM-based knowledge in the sentence embeddings through the stages while avoiding the cost of huge computation by training either audio or text encoder in a stage.
The experiments validate the effectiveness of M2D2 in achieving our goal of a general-purpose audio-language representation. The learned audio feature exhibits SOTA performance in fine-tuning, and the CLAP features demonstrate top performance in audio-language tasks. For a successful pre-training by multitasking M2D and CLAP, we find it essential to balance the loss between M2D and CLAP to control the powerful LLM-based semantic supervision. The audio projector capacity, LLM-based text encoder, and two-stage training strategy designs are also crucial, which we demonstrate in the ablation studies in Section IV-D.
@@
A general-purpose audio-language representation that provides CLAP features for tasks where audio-language alignment is beneficial and general-purpose audio features for other conventional audio tasks.
A new pre-training method, M2D2, that combines M2D and CLAP using LLM-based sentence embeddings and provides SOTA general-purpose audio features and top-performing CLAP features, achieving a general-purpose audio-language representation.
Empirical findings on combining SSL and CLAP in ablation studies and publicly available code with pre-trained weights111https://github.com/nttcslab/m2d for future studies.
The first study to show that an LLM-based sentence embedding model from the NLP domain helps learn effective audio and CLAP features.
@@ II-A Relationship with our previous work
This study redefines and extends our previous work, M2D-CLAP[11]. M2D2 differs from M2D-CLAP in improved audio and CLAP features achieved by the newly employed two-stage pre-training strategy, improved audio projector based on a transformer encoder, and the use of LLM-based sentence embeddings (Section III-B). To explore its effectiveness further, we also conducted experiments on audio-text retrieval (Section IV-C2), audio captioning (Section IV-C3), and ablation studies (Section IV-D) in addition to audio transfer learning (Section IV-B) and zero-shot classification (Section IV-C1).
To make the difference from the final M2D2 clear, we specifically denote M2D-CLAP as M2D20- in the experiments because it is a variant of M2D2-, the first-stage pre-trained M2D2.
@@ II-B Conventional audio representation models
Since the release of the large-scale dataset AudioSet[17], various supervised learning models such as PANNs[24], AST[25], and HTS-AT[26] that utilize class labels have followed. In contrast, SSL methods such as COLA[27] and BYOL-A[15] are general-purpose audio signal representations that do not depend on task supervision, such as labels. Masked prediction-based SSLs have progressed and demonstrated remarkable general-purpose performance. SSAST[28], MAE-AST[29], MSM-MAE[30], and Audio-MAE[31] learn by reconstructing the masked input, while M2D and ATST[32] learn by predicting the masked features. More advanced models include BEATs[33], which learn by predicting distilled and tokenized labels, and CED[34], which distills multiple large models. While they perform well on conventional audio tasks, their features do not align with languages.
@@ II-C CLAP models
Methods that take a similar approach to M2D2 are SLIP[36] in the image domain (combines SSL and CLIP), FLAP[8] (combines MAE[37] and CLAP), SupMAM-CLAP[38] (combines Audio-MAE and distillation of a CLAP model), Cacophony[10] (combines Audio-MAE, CLAP, and a captioning model in a two-stage training), and CoLLAT[9] (CLAP with frozen text encoders). Unlike these approaches, we combine M2D, CLAP, and the recent LLM-based sentence embeddings in a two-stage pre-training to learn a general-purpose audio-language representation that provides generalizable audio and CLAP features.
@@ III Proposed Method
We propose M2D2, which extends a masked prediction-based SSL method, M2D, to combine it with CLAP to learn general-purpose audio-language representations effective for a broad range of conventional audio and audio-language tasks.
@@ III-A2 Transfer learning
After the pre-training, only the audio encoder is transferred from the online network in downstream tasks. As the encoder accepts the input no longer than the predefined duration (e.g., 608 time steps) in the pre-training stage, the inference procedure handles variable-length audio inputs in two ways: one is by cutting the input with the predefined duration and repeatedly encoding the chunks; the other is by interpolating the positional encoding to the input audio duration (i.e., stretching the acceptable input duration). The latter is used for fine-tuning the encoder to a task with an input duration similar to the pre-training and the former is used in all other cases, such as linear evaluation.
@@ III-A3 M2D for X
III-B Proposed M2D2: M2D Second Generation
M2D2 is a second-generation M2D, an M2D-X variant that adds a CLAP training objective to implement a general-purpose audio-language representation. As shown in Fig. 2, it takes a multitask approach comprising (a) M2D and (b) CLAP that learns from paired audio and caption data. The CLAP extension has a semantic network branch that conducts contrastive learning among audio and text features to align them in a semantic feature space.
M2D2 employs a two-stage training strategy to learn the best of audio and CLAP features under a multitask of M2D and CLAP without the cost of huge computational resources. In the first stage, it learns effective audio features from M2D that also jointly learns from powerful LLM-based sentence embeddings by aligning the audio features with them through CLAP, aiming to achieve the best audio features that are also effective as CLAP audio features. In the second stage, it learns effective CLAP text features that align with the learned audio features in a semantic feature space. Through these stages, M2D2 learns effective audio and CLAP features.
@@ III-B1 First stage pre-training
In the first stage, the captions are encoded into embeddings in advance using the LLM-based text encoder, and the training process takes audio for M2D input while taking text embeddings for the corresponding caption. M2D then conducts its training process and feeds the audio encoder’s output feature to the semantic network. While the original M2D-X framework provides in addition to , M2D2 uses only for training stability. The audio projector converts the audio features into semantic features . The text embeddings are used as they are or mapped to the same dimension as using a learnable linear layer and handled as text semantic features . The training process does not involve the LLM-based sentence embedding model, which requires a large GPU memory.
This stage aims to train an effective audio encoder, so it is crucial that the training signal, semantic feature , precisely represents the semantics of the text. We use a powerful LLM-based sentence embedding model and distill its semantic knowledge into the audio encoder through CLAP. The audio projector is a transformer encoder to summarize so that the can better represent the semantics of an audio input.
@@ III-B2 Second stage pre-training
The second stage pre-trains the text encoder and audio projector using CLAP alone and freezes the audio encoder weights so that we preserve effective audio features learned in the first stage, as shown in Fig. 3. To facilitate the CLAP training using a precise loss, we make the audio feature for an audio input consistent by using a low masking ratio and inputting the entire audio sample as is.
@@
III-B3 M2D2 transfer learning
@@
III-B4 Additional pre-training stages
While the two training stages enable the learning of general-purpose features, the experiments show that these features still leave a gap with top performance in existing benchmarks if they are not fine-tuned. Therefore, to facilitate the comparison with previous studies, we also use two additional stages to bring the experimental conditions closer to those of existing studies.
Stage 1.1: AudioSet fine-tuning
Stage 2.1: WavCaps pre-training
This stage further pre-trains the second stage audio projector and text encoder to refine the CLAP features using the same dataset as in WavCaps[7] and with no masking of the audio input to learn the best text/audio alignment. We show that this process improves performance in audio-language tasks in Section IV-C.
III-B5 Training design and efficiency
Though other training strategies are possible, we pre-train the audio encoder first because we aim to exploit the knowledge from the LLM for both audio and text features. The M2D2 first stage enables that, unlike typical CLAP models that use a pre-trained audio encoder solely learned from audio data. To practically train a CLAP, we need pre-trained audio/text encoders, which is why we train the text encoder in the second stage. To enable effective training without a huge computing budget, we train encoders separately and use the LLM offline.
We utilize an audio projector only, unlike previous CLAP, which typically uses both audio and text projectors. An audio projector is crucial to bridge the gap between the descriptive general audio and abstract (semantic) CLAP features so that they can be learned in a single model. A text projector degrades performance in M2D2, as verified in Section IV-D3.
The staged pre-training allows for training under fewer computing resources, in addition to learning a well-performing representation. M2D2 learns from LLM features in the first stage, while the features are constant for the same captions. Therefore, M2D2 can substantially reduce resource requirements by pre-computing and retaining features for all captions and referring to pre-computed features instead of loading an LLM instance during training. In contrastive learning, a large batch size is critical for sufficient training. While FLAP, a similar method, requires GPUs to achieve a batch size of 2304, M2D2 can conduct training with 4 A100 GPUs (80GB) for a batch size of 2048.
@@ IV Experiments
To assess the effectiveness of M2D2 representations, we evaluated the audio features on conventional audio tasks under settings of frozen audio features (Sections IV-B1 and IV-B2) and fine-tuning audio features (Section IV-B3) and the CLAP features using three tasks: zero-shot classification (Section IV-C1), audio-text retrieval (Section IV-C2), and audio captioning (Section IV-C3).
We also studied how the pre-training design and parameters impact the final performance (Section IV-D), compared the CLAP features using visualizations (Section IV-D4), and show how M2D2 focuses on the audio patches to summarize CLAP audio features (Section IV-D4). In the following sections, we refer to the M2D2 models as follows:
M2D2-: First stage pre-trained representation. Combination of pre-trained audio encoder/audio projector and off-the-shelf LLM-based sentence embedding model as a text encoder.
M2D2: Second stage pre-trained representation. Combination of the first-stage audio encoder and the second-stage pre-trained audio projector/text encoder.
M2D2+: Second stage plus stage 2.1 pre-trained representation. Combination of the first-stage audio encoder and the stage 2.1 pre-trained audio projector/text encoder.
M2D2: Fine-tuned representation to the benchmarks. Combination of the audio encoder fine-tuned in stage 1.1 and the stage 2.1 pre-trained audio projector/text encoder.
M2D20-: The M2D-CLAP[11] model as an ablation variant of M2D2- for using a non-LLM text encoder and AudioSet only for pre-training audio data.

### v2

# M2D-CLAP: Exploring General-purpose Audio-Language Representations Beyond CLAP

@@ Abstract
Contrastive language-audio pre-training (CLAP), which learns audio-language representations by aligning audio and text in a common feature space, has become popular for solving audio tasks. However, CLAP’s audio features lack generalizability, whereas self-supervised learning (SSL) models offer general-purpose features that perform well across diverse audio tasks. We aim to develop a broadly applicable audio representation and hypothesize that a model that learns both general audio and CLAP features should achieve our goal, which we call a general-purpose audio-language representation. To implement our hypothesis, we propose M2D-CLAP, the first approach to jointly learn effective general audio and CLAP features. It extends an SSL masked modeling duo (M2D) by incorporating CLAP and utilizes LLM-based sentence embeddings. The training process consists of multiple stages. In the first stage, generalizable audio features are pre-trained via a multitask objective combining M2D and CLAP, with CLAP leveraging LLM-based semantic embeddings to distill semantic knowledge into them. In the following stages, CLAP features are pre-trained and refined with guidance from the learned audio features. Experiments demonstrated that M2D-CLAP learns high-performing general audio features (e.g., AudioSet mAP of 49.0, SOTA results in music tasks) and CLAP features, thereby enabling a general-purpose audio-language representation.
@@
Contrastive language-image pre-training (CLIP)[1] has enabled the alignment of two modalities and has significantly impacted diverse domains. In the audio domain, contrastive language-audio pre-training (CLAP) has extended CLIP to audio[2][3][4][5], followed by various advancements[6][7][8][9][10][11] that have exhibited superior performance for applications such as zero-shot classification and audio captioning[7][12][13]. CLAP can map audio and captions into features aligned in a common feature space; therefore, these features can be effective for various downstream application tasks since many task settings are typically related to language, such as inference of class labels.
While the CLAP models have exhibited versatility in audio-language tasks, their audio features do not generalize well in conventional audio tasks because they typically use audio encoders based on supervised learning models. The state-of-the-art (SOTA) audio representations are self-supervised learning (SSL) models known to generalize to diverse downstream tasks[14]. In contrast, the supervised learning models, and thus the CLAP models, suffer on tasks different from the pre-training, as we reported in [11][15]. For example, they perform well in ESC-50[16], an environmental sound classification similar to the pre-trained AudioSet[17] but poorly in VoxCeleb1[18], a speaker identification task. However, some studies have applied CLAP audio features to tasks where SSL audio features may perform well, such as [19][20].
@@
On the other hand, to learn the best of CLAP, a SOTA sentence embedding model based on a large-language model (LLM) would be a good candidate as a CLAP text encoder. In the natural language processing (NLP) domain, sentence embedding models have recently progressed to exploit an LLM[21]. LLMs have demonstrated overwhelming ability in language tasks[22] and have the potential to map audio captions into better semantic features based on knowledge accumulated in huge parameters learned from massive data.
To implement a general-purpose audio-language representation and achieve its highest potential, we propose M2D-CLAP, the first approach to simultaneously achieve effective general audio and CLAP features. It extends a masked prediction-based SSL, M2D[23], combines it with a CLAP objective, and exploits the strong LLM-based sentence embeddings, as shown in Fig. 1.
M2D-CLAP adopts a multi-stage pre-training strategy that leverages LLM-based sentence embeddings to learn high-quality features while avoiding substantial computational costs. In the first stage, M2D-CLAP performs multitask learning with M2D and CLAP objectives to acquire general-purpose audio features aligned with LLM-based sentence embeddings, aiming to distill LLM knowledge into audio representations. These features are then fine-tuned on AudioSet to further enhance their effectiveness, particularly for CLAP. In the subsequent stages, M2D-CLAP pre-trains a smaller non-LLM text encoder to learn CLAP text features that align with the learned audio features, aiming to transfer the knowledge into the text features. By training either the audio or text encoder at each stage, M2D-CLAP achieves efficient training without incurring the high computational cost.
The experiments validate the effectiveness of M2D-CLAP in achieving our goal of learning a general-purpose audio-language representation. The resulting audio and CLAP features show top performance both as frozen and fine-tuned audio features, as well as in audio-language tasks. For a successful pre-training by multitasking M2D and CLAP, we find it essential to balance the loss between M2D and CLAP to appropriately control the strong semantic supervision. Additionally, the audio projector capacity, the use of an LLM-based text encoder, and the multi-stage training strategy all prove to be critical design choices, as demonstrated in the ablation studies in Section IV-D.
@@
Introducing a general-purpose audio-language representation that provides general audio and CLAP features to serve a wide range of audio and audio-language tasks.
Proposing M2D-CLAP, the first (to our knowledge) method to jointly learn top-performing general-purpose audio and CLAP features, thereby realizing a general-purpose audio-language representation.
Providing empirical findings on combining SSL and CLAP, along with publicly available code and pre-trained weights111https://github.com/nttcslab/m2d to support future research.
The differences from our conference paper[11] are summarized in the next section.
@@ II-A Relationship with our previous work
We introduced M2D-CLAP in our previous work[11]. While it demonstrated the effectiveness, the approach left room for improvement, particularly in the performance of the CLAP features. This paper redefines and extends M2D-CLAP to enhance the effectiveness of its audio and CLAP features, leveraging a newly introduced multi-stage pre-training strategy, a transformer-based audio projector, and LLM-based sentence embeddings (Section III-B). To further assess its effectiveness, we also conducted experiments on audio-text retrieval, audio captioning, and ablation analyses in addition to transfer learning and zero-shot classification (Sections IV-B)–IV-D).
To make the difference from the redefined M2D-CLAP clear, we specifically denote it as M2D-CLAP and the previous M2D-CLAP[11] as M2D-CLAP. Note that M2D-CLAP is a variant of the first-stage M2D-CLAP.
@@ II-B Conventional audio representation models
Since the release of the large-scale dataset AudioSet[17], various supervised learning models such as PANNs[24], AST[25], and HTS-AT[26] that utilize class labels have followed. In contrast, SSL methods such as COLA[27] and BYOL-A[15] are general-purpose audio representations that do not depend on task supervision, such as labels. Masked prediction-based SSLs have progressed and demonstrated remarkable general-purpose performance. SSAST[28], MAE-AST[29], MSM-MAE[30], and Audio-MAE[31] learn by reconstructing the masked input, while M2D and ATST[32] learn by predicting the masked features. More advanced models include BEATs[33], which learn by predicting distilled and tokenized labels, and CED[34], which distills multiple large models. While they perform well on conventional audio tasks, their features do not align with languages.
@@ II-C CLAP models
Methods that take a similar approach to M2D-CLAP are SLIP[36] in the image domain (combines SSL and CLIP), FLAP[8] (combines MAE[37] and CLAP), SupMAM-CLAP[38] (combines Audio-MAE and distills a CLAP model), Cacophony[10] (combines Audio-MAE, CLAP, and a captioning model in a two-stage training), and CoLLAT[9] (CLAP with frozen text encoders). Unlike these approaches, we combine M2D, CLAP, and the recent LLM-based sentence embeddings in a multi-stage pre-training to learn a general-purpose audio-language representation that provides effective audio and CLAP features.
@@ III Proposed Method
We propose M2D-CLAP, which extends M2D to combine it with CLAP to learn general-purpose audio-language representations that provide useful features for a broad range of conventional audio and audio-language tasks.
@@ III-A2 Transfer learning
After the pre-training, only the audio encoder is transferred from the online network to the downstream tasks. As the encoder accepts the input no longer than the predefined duration (e.g., 608 time steps) in the pre-training stage, the inference procedure handles variable-length audio inputs in two ways: one is by cutting the input with the predefined duration and repeatedly encoding the chunks; the other is by interpolating the positional encoding to the input audio duration (i.e., stretching the acceptable input duration). The latter is used for fine-tuning the encoder to a task with an input duration similar to the pre-training, and the former is used in all other cases, such as linear evaluation.
@@ III-A3 M2D for X
III-B Proposed M2D-CLAP: M2D Meets CLAP
M2D-CLAP is an M2D-X variant that combines M2D and CLAP training objectives to implement a general-purpose audio-language representation. As shown in Fig. 2, it takes a multitask approach comprising (a) M2D and (b) CLAP that learns from paired audio and caption data. The CLAP extension has a semantic network branch that conducts contrastive learning among audio and text features to align them in a semantic feature space.
M2D-CLAP employs a multi-stage training strategy to learn the best of audio and CLAP features under a multitask of M2D and CLAP without the cost of huge computational resources. In the first stage, it learns effective audio features from M2D that also jointly learns from powerful LLM-based sentence embeddings by aligning the audio features with them through CLAP, aiming to achieve the best audio features that are also effective as CLAP audio features. In the second stage, it learns effective CLAP text features that align with the learned audio features in a semantic feature space. In each stage, it additionally fine-tunes features to achieve effective CLAP performance. Through these stages, M2D-CLAP learns effective audio and CLAP features.
@@ III-B1 First stage pre-training
In the first stage, captions are encoded in advance into embeddings using an LLM-based sentence embedding model, ensuring that the training process receives only audio inputs alongside the corresponding fixed text embeddings. M2D then conducts its training process and feeds the audio encoder’s output feature to the semantic network. While the original M2D-X framework provides in addition to , M2D-CLAP uses only for training stability. The audio projector converts the audio features into semantic features . The text embeddings are used as they are or mapped to the same dimension as using a learnable linear layer and handled as text semantic features . The training process does not involve the LLM-based sentence embedding model, which requires a large GPU memory.
This stage aims to train an effective audio encoder, so it is crucial that the training signal, semantic feature , precisely represents the semantics of the text. We use a powerful LLM-based sentence embedding model and distill its semantic knowledge into the audio encoder through CLAP. The audio projector is a transformer encoder to summarize so that the can better represent the semantics of an input audio.
@@
The first-stage pre-training combines supervised learning (CLAP) and SSL (M2D), enabling the audio encoder to learn effective features through their synergy. CLAP directs feature learning toward the semantics of the caption text but may miss aspects of sound not explicitly described. In contrast, SSL promotes general-purpose feature learning but lacks the guidance provided by textual supervision. By integrating both, the model learns semantically aligned yet generalizable representations. The learned audio features improve performance on tasks such as environmental sound and music compared to M2D alone, at the cost of slightly reduced performance on speech tasks, but with an overall performance gain. The experiments in Section IV-B demonstrate this trade-off.
Additional stage 1.1: AudioSet fine-tuning
@@ III-B2 Second stage pre-training
The second stage pre-trains the text encoder and audio projector using CLAP alone and freezes the audio encoder weights so that we preserve effective audio features learned in the first stage, as shown in Fig. 3. To facilitate CLAP training with a precise loss, we input the entire audio sample without truncation and apply a low masking ratio, allowing a nearly consistent loss calculation.
@@
Additional stage 2.1: WavCaps fine-tuning
This stage further fine-tunes the second-stage audio projector and text encoder to refine the CLAP features using the same dataset as in WavCaps[7] and with no masking of the audio input to learn the best text/audio alignment. We show that this process improves performance in audio-language tasks in Section IV-C.
III-B3 M2D-CLAP transfer learning
@@
III-B4 Training design and efficiency
While other strategies are possible, we pre-train the audio encoder first to leverage LLM knowledge for both audio and text features. The first stage of M2D-CLAP enables this, unlike typical CLAP models that use an existing audio encoder pre-trained only on audio data. To train a CLAP model in practice, pre-trained audio and text encoders are required; thus, we train the audio encoder first and the text encoder in the second stage. To keep training efficient without high computational cost, we train the encoders separately and use the LLM offline.
We utilize an audio projector only, unlike previous CLAP, which typically uses both audio and text projectors. While an audio projector is crucial to bridge the gap between the descriptive general audio and abstract (semantic) CLAP features, a text projector degrades performance in M2D-CLAP, as verified in Section IV-D3.
The staged pre-training not only enables learning a well-performing representation but also allows training with reduced computational resources. In the first stage, M2D-CLAP learns from LLM features that remain fixed for identical captions, enabling a substantial reduction in computational cost by pre-computing features for all captions. During training, these pre-computed features are used instead of loading the LLM instance, saving computational resources. In contrastive learning, a large batch size is critical for sufficient training. While FLAP, a similar method, requires GPUs to achieve a batch size of 2304, M2D-CLAP can conduct training with 4 A100 GPUs for a batch size of 2048.
@@ IV Experiments
To assess the effectiveness of M2D-CLAP representations, we evaluated the audio features on conventional audio tasks under settings of frozen audio features (Sections IV-B1 and IV-B2) and fine-tuning audio features (Section IV-B3) and the CLAP features using three tasks: zero-shot classification (Section IV-C1), audio-text retrieval (Section IV-C2), and audio captioning (Section IV-C3).
We also studied how the pre-training design and parameters impact the final performance (Section IV-D), compared the CLAP features using visualizations (Section IV-D4), and show how M2D-CLAP focuses on the audio patches to summarize CLAP audio features (Section IV-D4). The following sections refer to M2D-CLAP and its intermediate-stage variants used in the ablation studies as follows:
M2D-CLAP: Final M2D-CLAP representation pre-trained in all two plus two additional stages.
M2D-CLAP: First stage pre-trained representation. Combination of pre-trained audio encoder/audio projector and off-the-shelf LLM-based sentence embedding model as a text encoder.
M2D-CLAP: Second-stage pre-trained representation, combining the first-stage audio encoder with the second-stage audio projector/text encoder. The audio encoder is not fine-tuned on AudioSet; stage 1.1 is omitted.
M2D-CLAP: Second stage plus stage 2.1 pre-trained representation, combining the first-stage audio encoder and the stage 2.1 pre-trained audio projector/text encoder.
M2D-CLAP: The model from the initial study[11] as an ablation variant of M2D-CLAP for using a non-LLM text encoder and AudioSet only for audio data.